OpenStudy (anonymous):

A 5000-kg truck drives over a pothole in the road that causes thetruck to push down on all four springs a distance of 4.0 cm. If energy is conserved and the springs obey Hooke’s law, calculate thespring constant for each sprin

OpenStudy (nottim):

Steps to doing a physics question. 1. Draw a diagram. 2. List the variables. 3. Determine the equation(s) involved.

OpenStudy (nottim):

m=5000kg d= 4.0cm=0.04m You are looking for k=? hooke's law is f=-kx.

OpenStudy (anonymous):

Is the total energy = gravitational potential energy + elastic potential energy ?

OpenStudy (nottim):

i think so.

OpenStudy (nottim):

you're your own best thinker.

OpenStudy (anonymous):

How do you get force?

OpenStudy (s3a):

Force caused by gravity is: F = m(-g) = -mg = -(5000kg)(9.8N/kg) (negative means downward) F = -49000 N (negative means downward) F = k(Δy) -49000 = k(-4) (Δy is negative because the initial position is at 0 meters and the final position is at -4 metres so Δy = -4 - 0 = -4 metres. Also notice that Δy being negative means that the truck went downward by 4 meters when the pot hole was hit.) (When people say F = -kx, you treat x as the distance that was moved while ignoring the sign but the way I did it makes you systematically think about it.) -49000/(-4) = 49000/4 = k = 12250 N/m (This is the spring constant - it means that in order to move the spring by 1 meter from its natural position where it would remain if nothing was forcefully compressing or stretching it, 12250 N of force would be required.)

OpenStudy (anonymous):

it's actually 4cm.....

OpenStudy (anonymous):

but nice try tho. Your answer is well structured.

OpenStudy (s3a):

Lol, oops. Four meters would be quite a fall. :D